Cort Inc. was authorized to issue $1,000,000 of 8 percent bonds. On April 1, 2019, the corporation issued bonds with a face value of $200,000 at a price of 102.0. The bonds mature 10 years from the date of issue. Interest is payable semiannually on October 1 and April 1.


Using the data given above, what amount of premium will be amortized by Cort Inc. on October 1, 2019, using straight-line amortization? Prepare the general journal entry to record this amortization.

Complete this question by entering your answers in the tabs below.

  • Amortized Premium
  • General Journal

Using the data given above, what amount of premium will be amortized by Cort Inc. on October 1, 2019, using straight-line amortization?

Amortized premium on October 1, 2019

Journal entry worksheet

  • Record the amortization of premium for six months.

Note: Enter debits before credits.

Date General Journal Debit Credit
Oct 01, 2019
Answer #1

a) Issue price of bonds = 200000*1.02 = 204000

Premium on bonds payable = 4000

Amortized premium on october 1,2019 = 4000/20 = 200

b) Journal entry

Date General Journal Debit Credit
Oct 01, 2019 Interest expense 7800
Premium on bonds payable 200
Cash (200000*8%*6/12) 8000
(To record interest)